A particle of charge 3 x 10-9 C is located at point (0,0) in a certain coordinate system. Assuming all lengths are in meters, calculate the potential at the following points: A (3,3) and B (0,2). How much work is needed to take a particle of charge 2 x10-5 C from point A to point B?


1
Expert's answer
2022-03-07T12:21:39-0500

Answer

Potential at A and B

V=6.4V

V'=13.5

Now

work is needed to take a particle of charge 2 x10-5 C from point A to point B

Is given

"W=q(V-V') \\=2*10^{-5}*(6.4-13.5) \\=1.42*10^{-18}J"